OpenStudy (idealist10):

Solve the initial-value problem y'=(xy+y^2)/x^2, y(-1)=2.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (idealist10):

y=ux u=y/x y'=u+u'x y'=u+u^2 u'x=u^2 du/u^2=dx/x -1/u+ln abs(x)+C u=-1/(ln abs(x)+C) y=-x/(ln abs(x)+C) y(-1)=1/(0+C)=2 C=1/2 y=-2x/(ln abs(x)+1)

OpenStudy (anonymous):

Which phrase represents the algebraic expression 5x - 9

OpenStudy (idealist10):